# Semi-parametric bulk and tail regression using spline-based neural networks

### v1

@@ Abstract
Semi-parametric quantile regression (SPQR) is a flexible approach to density regression that learns a spline-based representation of conditional density functions using neural networks. As it makes no parametric assumptions about the underlying density, SPQR performs well for in-sample testing and interpolation. However, it can perform poorly when modelling heavy-tailed data or when asked to extrapolate beyond the range of observations, as it fails to satisfy any of the asymptotic guarantees provided by extreme value theory (EVT). To build semi-parametric density regression models that can be used for reliable tail extrapolation, we create the blended generalised Pareto (GP) distribution, which i) provides a model for the entire range of data and, via a smooth and continuous transition, ii) benefits from exact GP upper-tails without the need for intermediate threshold selection. We combine SPQR with our blended GP to create extremal semi-parametric quantile regression (xSPQR), which provides a flexible semi-parametric approach to density regression that is compliant with traditional EVT. We handle interpretability of xSPQR through the use of model-agnostic variable importance scores, which provide the relative importance of a covariate for separately determining the bulk and tail of the conditional density. The efficacy of xSPQR is illustrated on simulated data, and an application to U.S. wildfire burnt areas.
@@ 1 Introduction
Extreme quantile regression models provide a useful tool for risk assessment in a wide range of applied fields, such as econometrics (Chavez-Demoulin and Guillou, , 2018; Daouia et al., , 2019, 2023), insurance (Daouia et al., , 2024), natural hazard modelling (Richards and Huser, , 2022; Yadav et al., , 2023), chemometrics (Gardes and Stupfler, , 2019), flood risk assessment (Jóhannesson et al., , 2022; Pasche and Engelke, , 2024), and downscaling (Wang et al., , 2012). A common approach to extreme quantile regression follows via a parametric assumption on the upper-tails of the conditional distribution of a response variable $Y\in\mathbb{R}$ given a set of realised covariates $\mathbf{x}\in\mathbb{R}^{p}$ of a random vector $\mathbf{X}$. The de-facto standard assumption is that excesses of $Y$ above a high threshold $u(\mathbf{x})$ follow a generalised Pareto (GP) distribution (see, e.g., Coles, (2001) for discussion) with covariate-dependent shape and scale parameters, $\xi(\mathbf{x})\in\mathbb{R}$ and $\sigma_{u}(\mathbf{x})>0$, respectively; that is, one assumes the parametric GP regression model:
$(Y-u(\mathbf{x}))\>|\;(Y>u(\mathbf{x}),\mathbf{X}=\mathbf{x})\sim\mbox{GP}(\sigma_{u}(\mathbf{x}),\xi(\mathbf{x})),$
@@
where the scale $\sigma_{u}(\mathbf{x})$ is dependent on the threshold $u(\mathbf{x})$. A regression model of this form benefits from the asymptotic guarantees of the GP distribution (e.g., threshold stability; see Coles, , 2001), and allows for reliable estimation of extreme conditional $\tau$-quantiles (for $\tau\in(0,1)$ close to one); see Davison and Smith, (1990) for details.
@@
The choice of the intermediate threshold $u(\mathbf{x})$ is an ubiquitous problem in extreme value theory (EVT) (see, e.g., Murphy et al., , 2024). In a regression setting, the threshold $u(\mathbf{x})$ can be modelled as a function of covariates; choosing and estimating an appropriate model for $u(\mathbf{x})$ introduces uncertainty into the modelling procedure. Moreover, the additional complication that the scale $\sigma_{u}(\mathbf{x})$ depends on $u(\mathbf{x})$ makes interpretability of model estimates difficult (Richards et al., , 2023). To overcome the problem of threshold choice, one can replace the parametric GP model (for threshold excesses) with a model for the entire support of data that retains (exact or approximate) GP upper-tails, without requiring threshold estimation. Many such univariate models have been proposed in the literature (see review by Scarrott and MacDonald, , 2012). Examples include splicing or mixing of density functions (with at least one constituent density being the GP density; see, e.g., Frigessi et al., , 2002; Behrens et al., , 2004; de Melo Mendes and Lopes, , 2004; Carreau and Bengio, , 2007; Reynkens et al., , 2017; Castro-Camilo et al., , 2019), blending of distribution functions (Castro-Camilo et al., , 2022; Krakauer, , 2024), or parametric generalisations of the GP distribution (Papastathopoulos and Tawn, , 2013; Naveau et al., , 2016; Stein, 2021a, ; Stein, 2021b, ). Although these all-in-one models are compliant with traditional EVT, and can thus perform reliable extrapolation into the upper-tail, they often suffer in terms of flexibility as both the bulk and tail of the distribution are parametric. Bulk and tail models with non-parametric bodies have been proposed (e.g., Tancredi et al., , 2006; MacDonald et al., , 2011; Tencaliec et al., , 2020), but these are yet to be extended to the regression setting.
@@
After estimating $u(\mathbf{x})$, a further modelling choice for GP regression is the functional forms of $\sigma_{u}(\mathbf{x})$ and $\xi(\mathbf{x})$. Although classical GP regression models used linear models (e.g., Davison and Smith, , 1990; Wang et al., , 2012), there has been growing interest in more flexible and computationally-scalable alternatives, including additive models (Chavez-Demoulin and Davison, , 2005; Youngman, , 2019), local kernel smoothing (Daouia et al., , 2011; Velthoen et al., , 2019; Gardes and Stupfler, , 2019), decision trees (Farkas et al., , 2021, 2024), random forests (Gnecco et al., , 2024), gradient boosting (Koh, , 2023; Velthoen et al., , 2023), and deep neural networks (Wilson et al., , 2022; Richards et al., , 2023; Cisneros et al., , 2024; Pasche and Engelke, , 2024; Mackay et al., , 2024). In particular, deep parametric GP regression methods (i.e., those that exploit neural networks for modelling $\sigma_{u}(\mathbf{x})$ and $\xi(\mathbf{x})$; see overview by Richards and Huser, , 2024) are particularly appealing as they are highly flexible, scalable to high covariate dimension $p$, and can be estimated quickly using easily-accessible software packages (see, e.g., Richards, , 2022). However, existing deep GP regression models suffer from the same two limitations of all parametric GP regression models: i) they require modelling and estimation of the intermediate threshold $u(\mathbf{x})$, and ii) they do not provide a model for the bulk (or body) and lower-tails of the conditional distribution, that is, $Y\>|\;(Y<u(\mathbf{x}),\mathbf{X}=\mathbf{x})$.
@@
To create flexible, semi-parametric regression models that are EVT-compliant and capable of modelling the entire range of the response data, we introduce the blended GP (bGP) distribution. Similarly to the construction of the blended generalised extreme value distribution of Castro-Camilo et al., (2022), we blend the GP distribution with an appropriately chosen distributional model for the bulk of the data. The bGP distribution can model the entire range of data, and has exact GP upper-tails without the need for threshold selection. We utilise the blended GP for regression by combining a deep GP regression model, for the upper-tails, with a deep semi-parametric regression model for the bulk of the distribution. For the latter, we employ the semi-parametric quantile regression (SPQR) model of Xu and Reich, (2021), which has been used previously for inference with spatial extremal processes (Majumder and Reich, , 2023; Majumder et al., 2024b, ). Via a spline-based representation of conditional densities, the SPQR framework provides a flexible description of the conditional bulk without any parametric assumptions. While it has been shown to work extremely well for in-sample prediction, the SPQR model assumes finite support for the response variable $Y$, and so cannot be readily used to perform extrapolation beyond the range of the observations. However, in conjunction with a blended GP regression model, we create a very flexible density regression model that is EVT-compliant; we refer to this framework as extremal semi-parametric quantile regression (xSPQR).
@@
The contributions of this work are threefold: i) we introduce the (unconditional) blended GP (bGP) distribution, which provides a new univariate distribution that can jointly model the bulk and upper-tail of data; ii) we develop the xSPQR framework for all-in-one, EVT-compliant density regression and extrapolation; and iii) we showcase interpretability of the xSPQR method through the use of accumulative local effects (ALEs; Apley and Zhu, , 2020), which are a model-agnostic metric for assessing relative variable importance.
@@
The rest of this paper is organised as follows. In Section 2, we provide the background on SPQR. In Section 3, we introduce the blended GP distribution, and combine this with SPQR to create xSPQR, with details on estimation of variable importance provided in Section 3.5. In Section 4, we conduct simulation studies to demonstrate the efficacy of xSPQR and its ability to capture covariate effects on both the bulk and tails of the conditional density function. In Section 5, we apply xSPQR to model the entire range of burnt areas from moderate and extreme wildfire in the contiguous United States (U.S.) between 1990–2020. We conclude in Section 6 with a discussion, and outline avenues for further research.
@@ 2.2 Semi-parametric quantile regression
Semi-parametric quantile regression (SPQR) was introduced by Xu and Reich, (2021) as an approach to conditional density estimation which makes no parametric assumptions about the underlying distribution. Consider a response variable $Y\in[0,1]$ and a set of realised covariates $\mathbf{x}\in\mathcal{X}$ of a random vector X, where $\mathcal{X}\subset\mathbb{R}^{p}$ denotes the sample space of X. Then, SPQR assumes that the conditional density of $Y\mid\mbox{\bf X}=\mbox{\bf x}$ is a function of $K$ third-order $M$-spline basis functions, $\{M_{1}(y),...,M_{K}(y)\}$, with $K+2$ knots spanning the unit interval $[0,1]$. $M$-spline basis functions are supported on a compact interval, positive, and integrate to one, giving them a natural interpretation as density functions. Moreover, both the $M$-spline basis function and its integral, the $I$-spline basis function, have a closed-form construction, which permits efficient inference and prediction with SPQR; see Appendix A for details. We further note that a convex combination of $M$-spline ($I$-spline) basis functions is a valid density (distribution) function supported on a compact interval (Ramsay, , 1988). The SPQR model assumes that the conditional density can be represented as a convex combination of $M$-spline basis functions:
@@
By increasing the number $K$ of basis functions and appropriately selecting the weights, the mixture density in (1) can approximate any continuous density function (Chui et al., , 1980; Abrahamowicz et al., , 1992). To model the weights in a flexible manner, Xu and Reich, (2021) use a multi-layer perceptron (MLP), a type of neural network. The MLP has $H$ hidden layers with non-linear activation functions; common activation functions used in previous applications of SPQR include the sigmoid and rectified linear unit (ReLU, Nair and Hinton, , 2010) functions, which we consider in this work. The additional final layer of the MLP outputs the weights $\{w_{k}\}_{k=1}^{K}$ and has a softmax activation function, which ensures that the outputs sum to one. The width of each hidden layer is determined by the number of nodes, $n_{h}\in\mathbb{N}$, for $h=1,\dots,H$.
@@
$\mbox{\bf x}^{(h)}$ $:=\bm{\phi}(\mbox{\bf W}^{(h)}\mbox{\bf x}^{(h-1)}+\mbox{\bf b}^{(h)}),\quad h=1,\ldots,H,$
$\mathcal{W}(\mbox{\bf x}):=\mbox{\bf x}^{(H+1)}$ $=\mbox{softmax}(\mbox{\bf W}^{(H+1)}\mbox{\bf x}^{(H)}+\mbox{\bf b}^{(H+1)}),$ (2)
where $\bm{\phi}$ is the non-linear activation function, $^{(0)}:=\mbox{\bf x}$ x corresponds to the input covariates, and $\mathcal{W}(\mbox{\bf x})=\{w_{k}(\mbox{\bf x})\}_{k=1}^{K}$ is a $K$-vector of weights satisfying $\sum_{k=1}^{K}w_{k}(\mbox{\bf x})=1$ for all $\mbox{\bf x}\in\mathcal{X}$. The estimable “weights” and “biases” of the MLP are contained in $\bm{\theta}:=\{\mbox{\bf W}^{(h)},\mbox{\bf b}^{(h)}\}_{h=1}^{H+1}$, where $\mbox{\bf W}^{(h)}\in\mathbb{R}^{n_{h}\times n_{h-1}}$ and $\mbox{\bf b}^{(h)}\in\mathbb{R}^{n_{h}}$ are layer-specific weight matrices and bias vectors, respectively, with input dimension $n_{0}:=p$. Note that the dependence of $\mathcal{W}(\cdot)$ (and its elements) on $\bm{\theta}$ has been omitted from notation for simplicity. If each layer is represented by a function, say $f^{(h)}(\cdot)$ (suppressing the dependence on $\bm{\theta}$ and X for convenience), the MLP $f:\mathbb{R}^{q}\rightarrow\mathbb{R}^{K}$ is a composition of the layer operations, i.e., $f(\cdot):=f^{(H+1)}\circ\ldots\circ f^{(1)}(\cdot)$.
We use stochastic gradient descent and the adaptive moment estimation (Adam) optimizer (Kingma and Ba, , 2014) to optimise $\bm{\theta}$. The optimised loss function is the negative log-likelihood associated with (1). Hyper-parameters, such as the number of hidden layers $H$, the dimension $n_{h}$ of each hidden layer, the activation function $\bm{\phi}$, and the number of basis functions $K$, are tuned using a grid search. Computing the conditional density function (1) is simple and fast given the parameters $\bm{\theta}$. Additionally, the expression for the distribution function is also available in closed form:
$F_{\rm SPQR}(y|\mbox{\bf x})=\sum_{k=1}^{K}w_{k}(\mbox{\bf x})I_{k}(\mbox{\bf x}),$ (3)
where $\{I_{k}(\cdot)\}_{k=1}^{K}$ are $I$-spline basis functions; note that the weights $\{w_{k}(\mbox{\bf x})\}_{k=1}^{K}$ do not differ between the density and distribution function in (1) and (3), respectively. Finally, the conditional quantile function for level $\tau\in(0,1)$ is defined as $Q_{\rm SPQR}(\tau|\mbox{\bf x})$, such that
@@
While $Q_{\rm SPQR}(\tau|\mbox{\bf x})$ is not available in closed-form (unlike $f_{\rm SPQR}(y|\mbox{\bf x})$ and $F_{\rm SPQR}(y|\mbox{\bf x})$) it can be quickly approximated using interpolation. Finally, we note that SPQR can accommodate the special case where X is just an intercept term, which is equivalent to unconditional density estimation; this has been recently explored by Majumder et al., 2024a .

### v2

@@ Abstract
Semi-parametric quantile regression (SPQR) is a flexible approach to density regression that learns a spline-based representation of conditional density functions using neural networks. As it makes no parametric assumptions about the underlying density, SPQR performs well for in-sample testing and interpolation. However, it can perform poorly when modelling heavy-tailed data or when asked to extrapolate beyond the range of observations, as it fails to satisfy any of the asymptotic guarantees provided by extreme value theory (EVT). To build semi-parametric density regression models that can be used for reliable tail extrapolation, we create the blended generalised Pareto (GP) distribution, which i) provides a model for the entire range of data and, via a smooth and continuous transition, ii) benefits from exact GP upper-tails without the need for intermediate threshold selection. We combine SPQR with our blended GP to create semi-parametric quantile regression for extremes (SPQRx), which provides a flexible semi-parametric approach to density regression that is compliant with traditional EVT. We handle interpretability of SPQRx through the use of model-agnostic variable importance scores, which provide the relative importance of a covariate for separately determining the bulk and tail of the conditional density. The efficacy of SPQRx is illustrated on simulated data, and an application to U.S. wildfire burnt areas from 1990–2020.
@@ 1 Introduction
Extreme quantile regression models provide a useful tool for risk assessment in a wide range of applied fields, such as econometrics (Chavez-Demoulin and Guillou, 2018; Daouia et al., 2019, 2023), insurance (Daouia et al., 2024), natural hazard modelling (Richards and Huser, 2026b; Yadav et al., 2023), chemometrics (Gardes and Stupfler, 2019), flood risk assessment (Jóhannesson et al., 2022; Pasche and Engelke, 2024), and downscaling (Wang et al., 2012). A common approach to extreme quantile regression follows via a parametric assumption on the upper-tails of the conditional distribution of a response variable $Y\in\mathbb{R}$ given a set of realised covariates $\mathbf{x}\in\mathbb{R}^{p}$ of a random vector $\mathbf{X}$. The de-facto standard assumption is that excesses of $Y$ above a high threshold $u(\mathbf{x})$ follow a generalised Pareto (GP) distribution (see, e.g., Coles (2001) for discussion) with covariate-dependent shape and scale parameters, $\xi(\mathbf{x})\in\mathbb{R}$ and $\sigma_{u}(\mathbf{x})>0$, respectively. That is, one assumes the parametric GP regression model:
$\bigg(Y-u(\mathbf{x})\bigg)\>|\;\bigg(Y>u(\mathbf{x}),\mathbf{X}=\mathbf{x}\bigg)\sim\mbox{GP}\bigg(\sigma_{u}(\mathbf{x}),\xi(\mathbf{x})\bigg),$
@@
where the scale $\sigma_{u}(\mathbf{x})$ is dependent on the threshold $u(\mathbf{x})$. A regression model of this form benefits from the asymptotic guarantees of the GP distribution (e.g., threshold stability; see Coles, 2001), and allows for reliable estimation of extreme conditional $\tau$-quantiles (for $\tau\in(0,1)$ close to one); see Davison and Smith (1990) for details.
@@
The choice of the intermediate threshold $u(\mathbf{x})$ is an ubiquitous problem in extreme value theory (EVT) (see, e.g., Murphy et al., 2025). In a regression setting, the threshold $u(\mathbf{x})$ can be modelled as a function of covariates. Choosing and estimating an appropriate model for $u(\mathbf{x})$ introduces uncertainty into the modelling procedure. Moreover, the additional complication that the scale $\sigma_{u}(\mathbf{x})$ depends on $u(\mathbf{x})$ makes interpretability of model estimates difficult (Richards et al., 2023). To overcome the problem of threshold choice, one can replace the parametric GP model (for threshold excesses) with a model for the entire support of data that retains (exact or approximate) GP upper-tails, without requiring threshold estimation. Many such univariate models have been proposed in the literature (see review by Scarrott and MacDonald, 2012). Examples include splicing or mixing of density functions (with at least one constituent density being the GP density; see, e.g., Frigessi et al., 2002; Behrens et al., 2004; de Melo Mendes and Lopes, 2004; Carreau and Bengio, 2007; Reynkens et al., 2017; Castro-Camilo et al., 2019), blending of distribution functions (Castro-Camilo et al., 2022; Krakauer, 2024), or parametric generalisations of the GP distribution (Papastathopoulos and Tawn, 2013; Naveau et al., 2016; Stein, 2021a, b; de Carvalho et al., 2022). Although these all-in-one models are compliant with traditional EVT, and can thus perform reliable extrapolation into the upper-tail, they often suffer in terms of flexibility as both the bulk and tail of the distribution are parametric. Bulk and tail models with non-parametric bodies have been proposed (e.g., Tancredi et al., 2006; MacDonald et al., 2011; Tencaliec et al., 2020), but these are yet to be extended to the regression setting.
@@
After estimating $u(\mathbf{x})$, a further modelling choice for GP regression is the functional forms of $\sigma_{u}(\mathbf{x})$ and $\xi(\mathbf{x})$. Although classical GP regression models used linear models (e.g., Davison and Smith, 1990; Wang et al., 2012), there has been growing interest in more flexible and computationally-scalable alternatives, including additive models (Chavez-Demoulin and Davison, 2005; Youngman, 2019), local kernel smoothing (Daouia et al., 2011; Velthoen et al., 2019; Gardes and Stupfler, 2019), decision trees (Farkas et al., 2021, 2024), random forests (Gnecco et al., 2024), gradient boosting (Koh, 2023; Velthoen et al., 2023), and deep neural networks (Wilson et al., 2022; Richards et al., 2023; Cisneros et al., 2024; Pasche and Engelke, 2024; Mackay et al., 2025). In particular, deep parametric GP regression methods (i.e., those that exploit neural networks for modelling $\sigma_{u}(\mathbf{x})$ and $\xi(\mathbf{x})$; see overview by Richards and Huser (2026a), Ch. 21 of de Carvalho et al., 2026) are particularly appealing as they are highly flexible, scalable to high covariate dimension $p$, and can be estimated quickly using easily-accessible software packages (see, e.g., Richards, 2022). However, existing deep GP regression models suffer from the same two limitations of all parametric GP regression models: i) they require modelling and estimation of the intermediate threshold $u(\mathbf{x})$, and ii) they do not provide a model for the bulk (or body) and lower-tails of the conditional distribution, that is, $Y\>|\;(Y<u(\mathbf{x}),\mathbf{X}=\mathbf{x})$.
@@
To create flexible, semi-parametric regression models that are EVT-compliant and capable of modelling the entire range of the response data, we introduce the blended GP (bGP) distribution. Similarly to the construction of the blended generalised extreme value distribution of Castro-Camilo et al. (2022), we blend the GP distribution with an appropriately chosen distributional model for the bulk of the data. The bGP distribution can model the entire range of data, and has exact GP upper-tails without the need for threshold selection. We utilise the blended GP for regression by combining a deep GP regression model, for the upper-tails, with a deep semi-parametric regression model for the bulk of the distribution. For the latter, we employ the semi-parametric quantile regression (SPQR) model of Xu and Reich (2021), which has been used previously for inference with spatial extremal processes (Majumder and Reich, 2023; Majumder et al., 2024b). Via a spline-based representation of conditional densities, the SPQR framework provides a flexible description of the conditional bulk without any parametric assumptions. While it has been shown to work extremely well for in-sample prediction, the SPQR model assumes finite support for the response variable $Y$, and so cannot be readily used to perform extrapolation beyond the range of the observations. However, in conjunction with a blended GP regression model, we create a very flexible density regression model that is EVT-compliant. We refer to this framework as semi-parametric quantile regression for extremes (SPQRx).
@@
The contributions of this work are threefold: i) we introduce the (unconditional) blended GP (bGP) distribution, which provides a new univariate distribution that can jointly model the bulk and upper-tail of data; ii) we develop the SPQRx framework for all-in-one, EVT-compliant density regression and extrapolation; and iii) we showcase interpretability of the SPQRx method through the use of accumulative local effects (ALEs; Apley and Zhu, 2020), which are a model-agnostic metric for assessing relative variable importance.
@@
The rest of this paper is organised as follows. In Section 2, we provide the background on SPQR. In Section 3, we introduce the blended GP distribution, and combine this with SPQR to create SPQRx, with details on estimation of variable importance provided in Section 3.5. In Section 4, we conduct simulation studies to demonstrate the efficacy of SPQRx and its ability to capture covariate effects on both the bulk and tails of the conditional density function. In Section 5, we apply SPQRx to model the entire range of burnt areas from moderate and extreme wildfire in the contiguous United States (U.S.) between 1990–2020. We conclude in Section 6 with a discussion, and outline avenues for further research.
@@ 2.2 Semi-parametric quantile regression
Semi-parametric quantile regression (SPQR) was introduced by Xu and Reich (2021) as an approach to conditional density estimation that makes no parametric assumptions about the underlying distribution. Consider a response variable $Y\in[0,1]$ and a set of realised covariates $\mathbf{x}\in\mathcal{X}$ of a random vector X, where $\mathcal{X}\subset\mathbb{R}^{p}$ denotes the sample space of X. Then, SPQR assumes that the conditional density of $Y\mid\mbox{\bf X}=\mbox{\bf x}$ is a function of $K$ third-order $M$-spline basis functions, $\{M_{1}(y),...,M_{K}(y)\}$, with $K+3$ knots spanning the unit interval $[0,1]$. $M$-spline basis functions are supported on a compact interval, positive, and integrate to one, giving them a natural interpretation as density functions. Moreover, both the $M$-spline basis function and its integral, the $I$-spline basis function, have a closed-form construction, which permits efficient inference and prediction with SPQR; see Appendix A for details. We further note that a convex combination of $M$-spline (II-spline) basis functions is a valid density (distribution) function supported on a compact interval (Ramsay, 1988). The SPQR model assumes that the conditional density can be represented as a convex combination of $M$-spline basis functions:
@@
By increasing the number $K$ of basis functions and appropriately selecting the weights, the mixture density in (1) can approximate any continuous density function (Chui et al., 1980; Abrahamowicz et al., 1992). To model the weights in a flexible manner, Xu and Reich (2021) use a multi-layer perceptron (MLP), a type of neural network (for details, see, e.g., Bishop and Bishop, 2023). The MLP has $H$ hidden layers with non-linear activation functions. Common activation functions used in previous applications of SPQR include the sigmoid and rectified linear unit (ReLU, Nair and Hinton, 2010) functions, which we consider in this work. The additional final layer of the MLP outputs the weights $\{w_{k}\}_{k=1}^{K}$ and has a softmax activation function, which ensures that the outputs sum to one. The width of each hidden layer is determined by the number of nodes, $n_{h}\in\mathbb{N}$, for $h=1,\dots,H$.
@@
$\mbox{\bf x}^{(h)}$ $:=\bm{\phi}\left(\mbox{\bf W}^{(h)}\mbox{\bf x}^{(h-1)}+\mbox{\bf b}^{(h)}\right),\quad h=1,\ldots,H,$
$\mathcal{W}(\mbox{\bf x}):=\mbox{\bf x}^{(H+1)}$ $=\mbox{softmax}\left(\mbox{\bf W}^{(H+1)}\mbox{\bf x}^{(H)}+\mbox{\bf b}^{(H+1)}\right),$ (2)
where $\bm{\phi}$ is the non-linear activation function, $\mbox{\bf x}^{(0)}:=\mbox{\bf x}$ corresponds to the input covariates, and $\mathcal{W}(\mbox{\bf x})=\{w_{k}(\mbox{\bf x})\}_{k=1}^{K}$ is a $K$-vector of weights satisfying $\sum_{k=1}^{K}w_{k}(\mbox{\bf x})=1$ for all $\mbox{\bf x}\in\mathcal{X}$. The estimable “weights” and “biases” of the MLP are contained in $\bm{\theta}:=\{\mbox{\bf W}^{(h)},\mbox{\bf b}^{(h)}\}_{h=1}^{H+1}$, where $\mbox{\bf W}^{(h)}\in\mathbb{R}^{n_{h}\times n_{h-1}}$ and $\mbox{\bf b}^{(h)}\in\mathbb{R}^{n_{h}}$ are layer-specific weight matrices and bias vectors, respectively, with input dimension $n_{0}:=p$. Note that the dependence of $\mathcal{W}(\cdot)$ (and its elements) on $\bm{\theta}$ has been omitted from notation for simplicity. If each layer is represented by a function, say $f^{(h)}(\cdot)$ (suppressing the dependence on $\bm{\theta}$ and X for convenience), the MLP $f:\mathbb{R}^{q}\rightarrow\mathbb{R}^{K}$ is a composition of the layer operations, i.e., $f(\cdot):=f^{(H+1)}\circ\ldots\circ f^{(1)}(\cdot)$.
We use stochastic gradient descent and the adaptive moment estimation (Adam) optimizer (Kingma and Ba, 2014) to optimise $\bm{\theta}$. The optimised loss function is the negative log-likelihood associated with (1). Hyper-parameters, such as the number of hidden layers $H$, the dimension $n_{h}$ of each hidden layer, the activation function $\bm{\phi}$, and the number of basis functions $K$, are tuned using a grid search. Computing the conditional density function (1) is simple and fast given the parameters $\bm{\theta}$. Additionally, the expression for the distribution function is also available in closed form:
$F_{\rm SPQR}(y|\mbox{\bf x})=\sum_{k=1}^{K}w_{k}(\mbox{\bf x})I_{k}(y),$ (3)
where $\{I_{k}(\cdot)\}_{k=1}^{K}$ are $I$-spline basis functions. Note that the weights $\{w_{k}(\mbox{\bf x})\}_{k=1}^{K}$ do not differ between the density and distribution function in (1) and (3), respectively. Finally, the conditional quantile function for level $\tau\in(0,1)$ is defined as $Q_{\rm SPQR}(\tau|\mbox{\bf x})$, such that
@@
While $Q_{\rm SPQR}(\tau|\mbox{\bf x})$ is not available in closed-form (unlike $f_{\rm SPQR}(y|\mbox{\bf x})$ and $F_{\rm SPQR}(y|\mbox{\bf x})$) it can be quickly approximated using interpolation. Finally, we note that SPQR can accommodate the special case where X is just an intercept term, which is equivalent to unconditional density estimation; this has been recently explored by Majumder et al. (2024a).